OpenStudy (anonymous):

Determine the equation of the line tangent to the graph of the following function at x = 0. g(x) = ex(−4 + 3x + 3x2)

OpenStudy (anonymous):

the eqn is y+4-6x^2-3x

OpenStudy (anonymous):

y+4-6x^2-3x = 0 ***

OpenStudy (anonymous):

how did you figure that one man?

OpenStudy (anonymous):

first you the y-value when x=0(so y=-4) and then when you find the equation of the tangent at any piont (aka, the slope) you use the slope, and your x and y value and sub them into the y=mx+b eqn to get it or you can use slope=(-4-y)/(o-x)

OpenStudy (anonymous):

only one problem. we're looking for the equation of a straight line. so it couldn't have any squared values in it

OpenStudy (anonymous):

true, okay the answer is y=3x+4

OpenStudy (anonymous):

that was considered the wrong answer when i submitted

OpenStudy (anonymous):

what grade are you in?

OpenStudy (anonymous):

im a senior in high school

OpenStudy (anonymous):

okay, and what's the ex in front of the eqn?

OpenStudy (anonymous):

g(x) = e^x(−4 + 3x + 3x^2)

OpenStudy (anonymous):

e to the power of x?

OpenStudy (anonymous):

yezzir

OpenStudy (anonymous):

whoa thats wierd, i didnt even consider that

OpenStudy (anonymous):

just get the derivative using product rule, then sub 0 in for x and find the slope at that point, and then use y=mx+b, no big.

OpenStudy (anonymous):

a'ight could you check my answer for me after i work it?

OpenStudy (anonymous):

k go, young grasshopper

OpenStudy (anonymous):

is there answers in the back of your text book or something?